Sally paints a house in 4 hours, jill paints the same house in 6 hours, how long does it take them to paint the house together?: Sally paints a house in 4 hours, jill paints the same house in 6 hours, how long does it take them to paint the house together?
Answer by ptaylor(1326)

Let x= amount of time it takes them to paint the house together
Sally paints at the rate of 1/4 house per hour
Jill paints at the rate of 1/6 house per hour
Together they paint at the rate of 1/4 + 1/6 or 3/12 + 2/12=5/12 house per hour
So, our equation to solve:
(5/12)x=1 ( one house, that is) multiply each side by 12
5x=12 divide each side by 5
x=2 2/5 hours---------------------time it takes them to paint the house together

Hope this helps---ptaylor